# Flow-Aware Ellipsoidal Filtration for Persistent Homology of Recurrent Signals

### v1

@@ 6 Discussion
Across all experiments, adaptive ellipsoidal neighbourhoods captured the shape of the trajectory, even in data with changing frequencies and a narrow bottleneck. In the Hamiltonian system, they recovered the shape of attractor at smaller scales without joining the bottleneck too early, whereas Vietoris–Rips and Fermat filtrations connected dense regions prematurely and failed to recover the attractor shape (Fig. 3). In the synthetic chirp, they preserved both amplitude and signal shape through the bottleneck and gave the most accurate reconstruction, particularly for the small-amplitude part $y(t)$ (Figs. 4 and 5). On the 3D accelerometer data, they revealed one clear loop that remained stable across scales, while spherical averaging caused the structure to collapse at larger scales (Figs. 6 and 7).
These outcomes follow from how neighbourhoods are constructed. As described in Section 4, we assume that the point clouds are sampled from a flow. We then determine the local flow region as the union of spatial $k$-nearest neighbours and a short temporal window. PCA is applied to this union of points to estimate the variance in the local flow. By aligning ellipsoids with the principal components and stretching them according to the variances, we approximate the flow using ellipsoids as the scale increases. In contrast, the Vietoris–Rips and Fermat filtrations cannot incorporate this flow model.
In denoising (Section 5.2), all topological filters—where we examined the size of the neighbourhoods in terms of both cardinality ($k$-NN) and geometry (spherical and ellipsoidal) improved with SNR, but adaptive ellipsoids best preserved amplitude and phase in the bottleneck, yielding lower RMSE, especially for $y(t)$ (Figs. 4 and 5).
On 3D accelerometer signals (Section 5.3), ellipsoidal filtration maintained the loop structure across the four scales associated with the dominant $H_{1}$ class, whereas spherical averaging increasingly compressed the narrow passages (Figs. 6 and 7). Using the geometric median instead of the mean made the method more robust to outliers and revealed a clearer attractor structure (Fig. 8).
@@
For recurrence-time estimation (Section 5.4), scales derived from the most persistent $H_{1}$ class tie the recurrence threshold to the intrinsic geometry of the trajectory rather than to an arbitrary radius. Recurrence times were then obtained from the first-return intervals, as described in Section 2.1. Using these scales, the flow-aligned neighbourhoods detected first returns robustly under strong frequency drift and avoided spurious early returns across the bottleneck (Figs. 11 and 12).
@@
Since the dominant $H_{1}$ feature can shorten towards the diagonal or be replaced by spurious loops at low $\mathrm{SNR}$, selecting the denoising or recurrence threshold solely from the most persistent $H_{1}$ class may be unreliable. In our experiments (medium–high $\mathrm{SNR}$) with finite sample lengths, extreme excursions were uncommon and the dominant $H_{1}$ provided a stable guide; nevertheless, at lower $\mathrm{SNR}$ we caution against relying on a single $H_{1}$-based scale without additional checks.
Compared with the Vietoris–Rips filtration, the ellipsoidal approach introduces two additional parameters that control the number of spatial and temporal neighbours. The union of these neighbours defines the local region used to estimate the local flow geometry via PCA. The resulting eigenvectors indicate the dominant flow directions, while the eigenvalues determine how the ellipsoid is stretched along each axis. Although the ellipsoid is later scaled beyond this region during the filtration, its orientation and shape are derived from the local dynamics. This additional parametrisation improves adaptability but increases computational cost, since constructing the spatial tree and performing PCA for each point are computationally intensive steps.
In a broad sense, VR and ellipsoidal filtrations both consider entropy, in the context of uncertainty, as the scale parameter used to connect points. While VR creates isotropic (spherical) uncertainty regions around each point and increases these with the scale parameter, ellipsoidal filtration creates anisotropic (ellipsoidal) uncertainty regions. Notably, ellipsoidal filtration produces smaller ellipsoids in slow-flow regions and larger ones in fast-flow regions. In this way, it evaluates how the topology evolves over time without explicitly using time as a parameter but instead through entropy. When the uncertainty regions of two points intersect, we assume that, at that scale, they are no longer separable and therefore connect them to form edges and higher-dimensional simplices. An analogy can be drawn between ellipsoidal filtration and motion blur photography, as illustrated in Fig. 13
(a) Motion blur with temporal structure.
(b) Blur without temporal awareness.
@@
Figure 13: Analogy between ellipsoidal filtration and motion blur photography. (a) Motion blur with temporal structure (Image credit: Matti Sulanto, used with permission). (b) Blur without temporal awareness (Image credit: İnsunur Eryılmaz, used with permission).
7 Conclusion and Future Work
We propose a flow-aware ellipsoidal filtration that treats the observed point cloud as samples from a smooth flow. Using spatio–temporal neighbourhoods and local covariance (PCA), it fits ellipsoids aligned with the local direction and speed, yielding an adaptive anisotropic filtration. Future work includes comprehensive tests on real data sets and applications in higher-dimensional settings.

### v2

@@ VI Discussion
Across all experiments, adaptive ellipsoidal neighbourhoods preserved the trajectory geometry more consistently than the Vietoris–Rips and Fermat filtrations. In the Hamiltonian system, they recovered the structure at smaller scales without prematurely bridging the narrow bottleneck, whereas the alternative filtrations connected dense regions too early (Fig. 3). In the synthetic chirp, they retained both amplitude and waveform through the bottleneck and yielded the most accurate reconstruction, particularly for the small-amplitude component $y(t)$ (Figs. 4 and 5). For the three-dimensional accelerometer data, loop structure persisted across scales under ellipsoidal neighbourhoods, while spherical averaging collapsed the geometry at larger scales (Figs. 6 and 7).
This behaviour is consistent with an anisotropic neighbourhood model that adapts to local flow variability. By allowing the neighbourhood shape and orientation to vary with local geometry, the ellipsoidal construction can follow narrow passages without joining distinct regions too early. However, the current method does not explicitly use tangent vectors when selecting neighbours. Incorporating tangent-direction information into neighbour selection and/or complex construction may further improve noise robustness.
A complementary interpretation is to view the scale parameter as controlling the size of an uncertainty region around each sample. As scale increases, these regions expand and the filtration examines how separability changes: when two regions intersect, the corresponding samples are assumed to be no longer distinguishable at that resolution and are connected via edges and higher-dimensional simplices. The Vietoris–Rips filtration uses isotropic (spherical) uncertainty regions that grow uniformly with scale, whereas the ellipsoidal filtration uses anisotropic (ellipsoidal) regions whose shape is determined by local covariance and tangent-direction estimates from spatio–temporal neighbourhoods. In dynamical terms, a tangent (velocity) vector encodes change with respect to time, so shaping uncertainty regions using tangent information makes the neighbourhood model temporally informed, even though the filtration parameter remains a geometric scale. This contrast is analogous to spatial blur versus motion blur: isotropic growth resembles reducing spatial resolution, while anisotropic, direction-aware growth reflects temporal structure (Fig. 11).
For denoising (Section V.2), we varied neighbourhoods by both cardinality ($k$-nearest neighbours) and geometry (spherical versus ellipsoidal). Performance improved with increasing SNR for all topological filters, but ellipsoidal neighbourhoods better preserved amplitude and phase near the bottleneck and achieved lower RMSE, especially for $y(t)$ (Figs. 4 and 5). In contrast, for horizontally elongated data the Vietoris–Rips filtration could promote an incorrect hole to be the most persistent $H_{1}$, and increasing noise could inflate that persistence, making the resulting scale unreliable. In our experiments, the ellipsoidal construction was less affected because the added noise followed the data’s anisotropy, and the method adapts neighbourhood shape using local variance information. However, these topological filters remain effective only when the data retain a distinguishable structure; strong noise or outliers can obscure the geometry and lead to unreliable features.
@@
For recurrence-time estimation (Section V.4), using the most persistent $H_{1}$ class to select scales links the recurrence threshold to the geometry of the trajectory rather than to an arbitrary radius. Recurrence times were then computed from first-return intervals (Section II.1). With these scales, flow-aligned neighbourhoods detected first returns under strong frequency drift while avoiding early returns across the bottleneck (Figs. 10a and 10b). Temporal neighbours helped the neighbourhoods track changes in the non-autonomous flow. Additional noise made the $H_{1}$-based scale less reliable for Vietoris–Rips because $H_{1}$ persistence increased, whereas the ellipsoidal filtration remained more stable under the same conditions.
@@
A key limitation is sensitivity to outliers: both Vietoris–Rips and ellipsoidal filtrations can produce spurious loops at low $\mathrm{SNR}$. The real three-dimensional accelerometer data (Fig. 7) illustrate a noise regime in which the dominant $H_{1}$ can guide denoising or recurrence analysis. In our high-SNR experiments with finite sample lengths, extreme excursions were uncommon and the dominant $H_{1}$ provided a stable guide. At lower SNR, persistence can shrink towards the diagonal or be replaced by spurious loops, so selecting a denoising or recurrence threshold from the most persistent $H_{1}$ should be treated with caution and complemented by additional checks.
Another limitation concerns the estimation of tangent vectors for constructing ellipsoids in higher-dimensions. As the ambient dimension increases, the number of samples needed for stable estimation grows rapidly, rendering the present implementation impractical in high dimensions. Moreover, the estimated tangent directions depend on the choice of neighbourhood size and temporal window, so different parameter choices may yield different ellipsoid orientations and reduce reproducibility across datasets. Diffusion‐geometry methods offer alternative strategies for local tangent estimation without assuming a global manifold and without reliance on explicit neighbourhood parameters, and integrating such techniques may improve robustness while preserving the geometric intent of the proposed construction (Jones, 2026).
Several directions remain for future work. Incorporating temporal information into diffusion geometry methods might increase the robustness of tangent vector estimation and reduce the number of samples required. Additionally, incorporating tangent-direction information into neighbour selection and the construction of simplicial complexes could enhance robustness. Robust persistent homology methods could be adapted for denoising (Anai et al., 2020; Damrich et al., ) and may help mitigate sensitivity to outliers. Finally, persistent-homology-guided scale selection could be explored using alternative similarity measures from recurrence analysis to derive more principled recurrence thresholds (Marwan and Kraemer, 2023).
@@
Figure 11: Analogy between ellipsoidal filtration and motion blur photography. Left: motion blur with directional structure. Right: blur without directional structure. Image credits: Matti Sulanto (left) and İnsunur Eryılmaz (right), used with permission.
VII Conclusion
We propose a flow-aware ellipsoidal filtration that treats the observed point cloud as samples from a smooth flow. Using spatio–temporal neighbourhoods and local covariance (PCA), the method fits locally aligned ellipsoids, yielding an adaptive anisotropic filtration. Across our experiments, this construction preserved trajectory geometry through bottlenecks more reliably than isotropic alternatives and provided stable, geometry-informed scales for denoising and recurrence-time estimation.
Acknowledgements.
O. B. Eryilmaz acknowledges funding from the Ministry of National Education of the Republic of Türkiye.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Appendix A Additional accelerometer denoising results
In this appendix we present the accelerometer denoising results obtained using geometric median averaging, complementing the mean-averaging results shown in Fig. 7.
[…]